# Automated Model Tuning for Multifidelity Uncertainty Propagation in Trajectory Simulation

### v1

@@ 5 Discussion and Conclusions
The practicality of automated model tuning for multifidelity UQ was explored in the context of an EDL application. When considered an offline cost, expending computational resources on tuning low-fidelity models to perform better with multifidelity estimators is a largely worthwhile exercise. However, for online analyses, such as those required for onboard trajectory simulation, the cost vs. precision tradeoff posed by model tuning optimization must be carefully considered.
@@
To explore this tradeoff, this study focused on implementing online multifidelity estimators using model tuning for a budget-constrained trajectory simulation problem. The time steps of two low-fidelity flight mechanics models were considered as tunable hyperparameters. A bi-level optimization strategy was employed to solve the model tuning problem, with the EGO algorithm on the outer loop and the ACV sample allocation problem on the inner loop, both implemented using Sandia National Laboratories Dakota software. The performance of using low-fidelity models that were tuned on-the-fly as part of an online multifidelity UQ analysis was assessed in terms of their variance and mean squared error with respect to three baseline approaches: 1) standard Monte Carlo, 2) ACVs with hand-tuned low-fidelity models, and 3) ACVs with globally optimal low-fidelity models, representing the best-case performance. The comparison was carried out for varying values of parameters that control the accuracy and cost of the model tuning optimization problem, namely the number of EGO iterations and the number of pilot samples used in each iteration to solve the inner loop ACV problem. Three trajectory simulation quantities of interest were considered across a range of total computational budgets.
@@
The results confirmed the intuition that the feasibility and usefulness of automated model tuning, when considering the overhead incurred, increases with the overall computational budget afforded for multifidelity UQ. For the particular EDL application studied, it was found incorporating model tuning largely outperforms multifidelity UQ with hand-tuned models, even when the total computational budget is relatively low. As the computational budget grows, estimators using model tuning approach the best-case performance possible for multifidelity UQ. Moreover, it was shown that incorporating a parametric ACV search over permissible model recursion graphs within the model tuning optimization could increase the potential variance reduction significantly. It is also worth noting that the viability of hand-tuning of hyperparameters is questionable in the case of on-board trajectory simulation, where environments and trajectories might be changing significantly from nominal values. In these cases, despite the small computational budget that may be afforded, on-the-fly model tuning is likely the best option.
@@
While this study sheds light on the utility of model tuning in a realistic setting, there remain open questions and avenues for improving the viability of such approaches. For example, the findings reported here with only two tunable low-fidelity models likely do not reflect the challenges that would be encountered in higher-dimensional model tuning problems. When several tunable models are present, the resulting optimization problem gets more complex, but the potential performance improvements likely grow significantly as hand-tuning becomes infeasible. In terms of future avenues for algorithmic improvements, developing cost-aware optimization algorithms for solving model tuning would allow the model tuning cost to be prescribed explicitly, rather than approximately through associated parameters (number of iterations and pilot samples). Such flexibility would help navigate the tradeoff of estimator cost versus precision with more rigor. Alternatively, a more dynamic assessment of diminishing returns from model tuning could be used to terminate the model tuning optimization problem, rather than the fixed convergence criteria considered here based on maximum number of iterations.
@@ Appendix B Example 1: Detailed Model Tuning Optimization Results
Detailed model tuning results for example 1 are shown in Figures 16 and 17 for varying numbers of pilot samples and EGO iterations, respectively. The top figures are scatter plots that show the estimated variance for each of the optimized time steps over 100 random trials. There is significant variability in the estimated variances, estimated with $N_{pilot}\leq 100$, versus the reference variance curve that was estimated with $N_{pilot}=50k$. While Figure 16 shows the relative scatter in both the estimated variance and optimized time step decrease with increasing $N_{pilot}$, this variability is relatively unchanged with increasing $N_{iter}$ in Figure 17
Figure 16: Model tuning optimization results for varying numbers of pilot samples with five EGO iterations across 100 random trials. The figures show the estimated optimal time step and estimator variance with the circular markers versus the reference variance profile (top) along with a histogram of the optimal time step values found by EGO (bottom).
Figure 17: Model tuning optimization results for varying numbers of EGO iterations with 25 pilot samples across 100 random trials. The figures show the estimated optimal time step and estimator variance with the circular markers versus the reference variance profile (top) along with a histogram of the optimal time step values found by EGO (bottom).
@@ Appendix C Model Tuning Optimization Problem Noise Illustration
This section helps illustrate the challenge posed by the noisy objective function for model tuning optimization for the ADEPT application. Figure 18 shows how estimates of the estimator variance (the objective function) vary across random trials for different numbers of pilot samples. Here, the estimator variance was evaluated on a grid of $25$ evenly spaced values of $log(\Delta t)$ for the time of flight QoI considered in Example 1. This evaluation was performed for $50$ random trials using $N_{pilot}=\{10,25,50,100\}$ and compared to the reference estimator variance curve computed with $N_{pilot}=50k$. Additionally, the optimal time step was found as the minimum across the $25$ $log(\Delta t)$ values considered for each trial. The figure shows significant variability in the vertical location of the estimator variance curves as well as the optimal time steps found (shown in the histograms in the lower plots). While the variability decreases as $N_{pilot}$ is increased, there is still a relatively wide range of optimal time step values for $N_{pilot}=100$. This illustrates that even for ideal optimizer performance (using a brute force sweep), it remains a challenge to optimize the estimator variance objective due to the noise associated with the ACV optimization problem and finite pilot sampling used.
Figure 18: Illustration of the noise in the model tuning optimization problem. The figures show the variability in estimator variance versus $log(\Delta t)$ across 50 random trials (top) along with the optimal value for $log(\Delta t)$ in each trial (bottom) when using a) 10, b) 25, c) 50, and d) 100 pilot samples to estimate model statistics.

### v2

@@ 5 Discussion and Conclusions
The practicality of automated model tuning for multifidelity uncertainty propagation was explored in the context of an EDL application. When considered an offline cost, expending computational resources on tuning low-fidelity models to perform better with multifidelity estimators is a largely worthwhile exercise. However, for online analyses, such as those required for onboard trajectory simulation, the cost vs. precision tradeoff posed by model tuning optimization must be carefully considered.
@@
To explore this tradeoff, this study focused on implementing online multifidelity estimators using model tuning for a budget-constrained trajectory simulation problem. The time steps of two low-fidelity flight mechanics models were considered as tunable hyperparameters. A bi-level optimization strategy was employed to solve the model tuning problem, with the EGO algorithm on the outer loop and the ACV sample allocation problem on the inner loop, both implemented using Sandia National Laboratories Dakota software. The performance of using low-fidelity models that were tuned on-the-fly as part of an online multifidelity uncertainty propagation analysis was assessed in terms of their variance and mean squared error with respect to three baseline approaches: 1) standard Monte Carlo, 2) ACVs with hand-tuned low-fidelity models, and 3) ACVs with globally optimal low-fidelity models, representing the best-case performance. The comparison was carried out for varying values of parameters that control the accuracy and cost of the model tuning optimization problem, namely the number of EGO iterations and the number of pilot samples used in each iteration to solve the inner loop ACV problem. Three trajectory simulation quantities of interest were considered across a range of total computational budgets.
@@
The results confirmed the intuition that the feasibility and usefulness of automated model tuning, when considering the overhead incurred, increases with the overall computational budget afforded for multifidelity uncertainty propagation with ACV estimators. For the particular EDL application studied, it was found incorporating model tuning largely outperforms ACV estimators with hand-tuned models, even when the total computational budget is relatively low. As the computational budget grows, estimators using model tuning approach the best-case ACV performance possible when optimal hyperparameters and model statistics are known a priori. Moreover, it was shown that incorporating a parametric ACV search over permissible model recursion graphs within the model tuning optimization could increase the potential variance reduction significantly. It is also worth noting that the viability of hand-tuning of hyperparameters is questionable in the case of on-board trajectory simulation, where environments and trajectories might be changing significantly from nominal values. In these cases, despite the small computational budget that may be afforded, on-the-fly model tuning is likely the best option.
@@
While this study sheds light on the utility of model tuning in a realistic setting, there remain open questions and avenues for improving the viability of such approaches. For example, the findings reported here with only two tunable low-fidelity models likely do not reflect the challenges that would be encountered in higher-dimensional model tuning problems. When several tunable models are present, the resulting optimization problem gets more complex, but the potential performance improvements likely grow significantly as hand-tuning becomes infeasible. In terms of future avenues for algorithmic improvements, employing advanced Bayesian optimization algorithms for solving model tuning such as those that are cost-aware [43] or more robust to noise [44] would help navigate the tradeoff of estimator cost versus precision with more rigor. Alternatively, a more dynamic assessment of diminishing returns from model tuning could be used to terminate the model tuning optimization problem, rather than the fixed convergence criteria considered here based on maximum number of iterations. Finally, this work adopted online pilot (iterated ACV) [33] as a simple method to handle unknown model statistics for multifidelity estimation, but more advanced adaptive approaches from recent work [45, 46, 47] could further improve estimator performance.
@@ Acknowledgements
James Warner and Geoffrey Bomarito carried out this work under the Entry Systems Modeling Project, which was funded by NASA’s Game Changing Development Program within the Space Technology Mission Directorate.
@@ Appendix B Example 1: Detailed Model Tuning Optimization Results
Detailed model tuning results for example 1 are shown in Figures 16 and 17 for varying numbers of pilot samples and EGO iterations, respectively. The top figures are scatter plots that show the estimated variance for each of the optimized time steps over 100 random trials. There is significant variability in the estimated variances, estimated with $N_{pilot}\leq 100$, versus the reference variance curve that was estimated with $N_{pilot}=10k$. While Figure 16 shows the relative scatter in both the estimated variance and optimized time step decrease with increasing $N_{pilot}$, this variability is relatively unchanged with increasing $N_{iter}$ in Figure 17
Figure 16: Model tuning optimization results for varying numbers of pilot samples with circular markers indicating the estimated optimal time step and estimator variance.
Figure 17: Model tuning optimization results for varying numbers of EGO iterations with circular markers indicating the estimated optimal time step and estimator variance.
@@ Appendix C Model Tuning Optimization Problem Noise Illustration
This section helps illustrate the challenge posed by the noisy objective function for model tuning optimization for the ADEPT application. Figure 18 shows how estimates of the estimator variance (the objective function) vary across random trials for different numbers of pilot samples. Here, the estimator variance was evaluated on a grid of $25$ evenly spaced values of $log(\Delta t)$ for the time of flight QoI considered in Example 1. This evaluation was performed for $50$ random trials using $N_{pilot}=\{10,25,50,100\}$ and compared to the reference estimator variance curve computed with $N_{pilot}=10k$. Additionally, the optimal time step was found as the minimum across the $25$ $log(\Delta t)$ values considered for each trial. The figure shows significant variability in the vertical location of the estimator variance curves as well as the optimal time steps found (shown in the histograms in the lower plots). While the variability decreases as $N_{pilot}$ is increased, there is still a relatively wide range of optimal time step values for $N_{pilot}=100$. This illustrates that even for ideal optimizer performance (using a brute force sweep), it remains a challenge to optimize the estimator variance objective due to the noise associated with the ACV optimization problem and finite pilot sampling used.
Figure 18: Variability in estimator variance versus $log(\Delta t)$ across 50 random trials when using a) 10, b) 25, c) 50, and d) 100 pilot samples.